# Asymptotic theory of in-context learning by linear attention

### v1

@@ 2.5 Model pretraining
We estimate model parameters using ridge regression, giving
@@
where $\lambda>0$ is a regularization parameter, and $H_{Z^{\mu}}$ refers to the input matrix (16) populated with the $\mu$th sample sequence. The factor $n/d$ in front of $\lambda$ makes sure that, when we take the $d\to\infty$ or $n\to\infty$ limits later, there is still a meaningful ridge regularization. The solution to the optimization problem in (19) can be expressed explicitly as
$\text{vec}(\Gamma^{\ast})=\left[\frac{n}{d}\lambda I+\sum_{\mu=1}^{n}\text{vec}(H_{Z^{\mu}})\text{vec}(H_{Z^{\mu}})^{\top}\right]^{-1}\sum_{\mu=1}^{n}y_{\ell+1}^{\mu}\text{vec}(H_{Z^{\mu}}),$ (20)
@@ 2.6 Evaluation
$e(\Gamma)=\mathbb{E}_{\mathcal{P}_{\mathrm{test}}}\left[\left(y_{\ell+1}-\langle\Gamma,H_{Z}\rangle\right)^{2}\right],$ (21)
where $(Z,y_{\ell+1})\sim\mathcal{P}_{\mathrm{test}}$ is a new sample drawn from the distribution of the test data set. We consider two different test data distributions $\mathcal{P}_{\mathrm{test}}$:
@@
    ICL task: $x_{i}$ and $\epsilon_{i}$ are i.i.d. Gaussians as in the pretraining case. However, for each $1\leq\mu\leq n$, the task vector $w^{\mu}$ associated with the $\mu$th input sequence is drawn independently from $\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$. We will denote the test error under this setting by $e^{\mathrm{ICL}}(\Gamma)$.
@@
    In-distribution generalization (IDG) task: The test data are generated in exactly the same manner as the training data, i.e., $\mathcal{P}_{\mathrm{test}}=\mathcal{P}_{\mathrm{train}}$, hence the term in-distribution generalization. In particular, the set of unique task vectors $\{w_{1},\ldots,w_{k}\}$ is identical to that used in the pretraining data. We will denote the test error under this setting by $e^{\text{IDG}}(\Gamma)$.
The ICL task evaluates the true in-context learning performance of the linear attention module. The task vectors in the test set differ from those seen in training, requiring the model to infer them from context. The IDG task assesses the model’s performance on task vectors encountered during pretraining. High performance on the IDG task but low performance on the ICL task indicates that the model memorizes the training task vectors. Conversely, high performance on the ICL task indicates that the model can learn task vectors from the provided context.
To understand the performance of our model on both ICL and IDG tasks, we will need to evaluate these expressions for the pretrained attention matrix $\Gamma^{\ast}.$ An asymptotically precise prediction of $e^{\mathrm{ICL}}(\Gamma^{\ast})$ and $e^{\mathrm{IDG}}(\Gamma^{\ast})$ will be a main result of this work.
2.7 Bayes optimal estimators
Following Raventós et al. [18], it is useful to compare the predictions made by the trained linear attention to optimal estimators that use only the current context information. These estimators do not rely on data outside of the given context for their predictions. Under the mean square loss, the optimal Bayesian estimator $\hat{y}_{\text{Bayes}}=\mathbb{E}_{\mathcal{P}_{\mathrm{test}}}[y_{\ell+1}|x_{1},y_{1},x_{2},y_{2},\ldots,x_{\ell},y_{\ell},x_{\ell+1}]$ in our setting has the form
$\hat{y}_{\text{Bayes}}=(w^{\text{Bayes}})^{\top}x_{\ell+1},$ (22)
where $w^{\text{Bayes}}$ is the Bayes estimator of the task vector $w$.
@@ 2.7 Bayes optimal estimators
$w^{\text{Bayes}}_{\text{ridge}}\coloneqq\left(\sum_{i=1}^{\ell}x_{i}x_{i}^{\top}+\rho I_{d}\right)^{-1}\left(\sum_{i=1}^{\ell}y_{i}x_{i}\right),$ (23)
where the ridge is set to the noise variance $\rho$. We will refer to it as the ridge estimator. For the IDG task, the Bayes-optimal estimator is given by
$w^{\text{Bayes}}_{\text{dMMSE}}\coloneqq\frac{\sum_{j=1}^{k}w_{j}e^{-\frac{1}{2\rho}\sum_{i=1}^{\ell}\left(y_{i}-w_{j}^{\top}x_{i}\right)^{2}}}{\sum_{j=1}^{k}e^{-\frac{1}{2\rho}\sum_{i=1}^{\ell}\left(y_{i}-w_{j}^{\top}x_{i}\right)^{2}}}.$ (24)
@@
$e^{\text{Bayes}}=\mathbb{E}_{\mathcal{P}_{\mathrm{test}}}\left[\left(y_{\ell+1}-(w^{\text{Bayes}})^{\top}x_{\ell+1}\right)^{2}\right],$ (25)
where $\mathcal{P}_{\mathrm{test}}$ can be the ICL or IDG task, and $w^{\text{Bayes}}$ can be the ridge or the dMMSE estimator. To avoid possible confusion, we emphasize that we will sometimes plot the performance of an estimator on a task for which it is not optimal. For example, we will test the dMMSE estimator, which is Bayes-optimal for the pretraining distribution, on the ICL task, where it is not optimal. This will be done for benchmarking purposes.
3 Theoretical results
To answer the questions raised in the introduction, we provide a precise asymptotic analysis of the learning curves of the linear attention module for ICL of linear regression. We then verify through simulations that the primary insights gained from our theoretical analysis extend to more realistic nonlinear Transformers.
3.1 Joint asymptotic limit
We have now defined both the structure of the training data as well as the parameters to be optimized. For our theoretical analysis, we consider a joint asymptotic limit in which the input dimension $d$, the pretraining dataset size $n$, the context length $\ell$, and number of task vectors in the training set $k$, go to infinity together such that
@@ 3.1 Joint asymptotic limit
The intuition for these scaling parameters can be seen as follows. Standard results in linear regression [34, 35, 36] show that to estimate a $d$-dimensional task vector $w$ from the $\ell$ samples within a context, one needs at least $\ell=\Theta(d)$. The number of unique task vectors that must be seen to estimate the true $d$-dimensional Gaussian task distribution should also scale with $d$, i.e. $k=\Theta(d)$. Finally, we see from (15) that the number of linear attention parameters to be learned is $\Theta(d^{2})$. This suggests that the number of individual contexts the model sees during pretraining should scale similarly, i.e., $n=\Theta(d^{2})$.
3.2 Learning curves for ICL of linear regression by a linear attention module
Our theoretical analysis, explained in detail in the Supplementary Information, leads to asymptotically precise expressions for the generalization error under the two test distributions under study. Specifically, our theory predicts that, as $d,n,\ell,k\to\infty$ in the joint limit given in (26),
@@ 3.2 Learning curves for ICL of linear regression by a linear attention module
where $e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$ and $e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)$ are two deterministic functions of the parameters $\tau$, $\alpha$, $\kappa$, $\rho$ and $\lambda$. The exact expressions of these two functions can be found in SI.5.2 and SI.5.3, respectively. For simplicity, we only present in what follows the ridgeless limit (i.e., $\lambda\to 0^{+}$) of the asymptotic generalization errors.
@@
$q^{\ast}\coloneqq\frac{1+\rho}{\alpha},\qquad m^{\ast}\coloneqq\mathcal{M}_{\kappa}\left({q^{\ast}}\right),\qquad\text{and}\qquad\mu^{\ast}\coloneqq q^{\ast}\mathcal{M}_{\kappa/\tau}(q^{\ast}),$ (29)
where $\mathcal{M}_{\kappa}(\cdot)$, defined in (B.3), is a function related to the Stieltjes transform of the Marchenko-Pastur law. Then
@@
where $\mathcal{M}^{\prime}_{\kappa}(\cdot)$ denotes the derivative of $\mathcal{M}_{\kappa}(q)$ with respect to $q$.
@@
$e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ $\coloneqq\lim_{\lambda\to 0^{+}}e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)$ (31)
=$=\begin{cases}\frac{\tau}{1-\tau}\left(\frac{\rho+q^{\ast}-2q^{\ast}(1-\tau)({q^{\ast}}/{\xi^{\ast}}+1)}{1-p^{\ast}(1-\tau)}+\frac{\tau\mu^{\ast}(q^{\ast}+\xi^{\ast})^{2}}{q^{\ast}}\right)&\qquad\tau<1\\ \frac{\tau}{\tau-1}[\rho+q^{\ast}(1-q^{\ast}m^{\ast})]&\qquad\tau>1\end{cases},$ (32)
@@
We derive these results using techniques from random matrix theory. The full setup and technical details are presented in the Supplementary Information in SI.1 through SI.5. A key technical component of our analysis involves characterizing the spectral properties of the sample covariance matrix of $n=\Theta(d^{2})$ i.i.d. random vectors in dimension $\Theta(d^{2})$. Each of these vectors is constructed as the vectorized version of the matrix in (16). Related but simpler versions of this type of random matrices involving the tensor product of i.i.d. random vectors have been studied in recent work [37]. Some of our derivations are based on non-rigorous yet technically plausible heuristics. We support these predictions with numerical simulations and discuss in the Supplementary Information the steps required to achieve a fully rigorous proof.
Before discussing the implications of our theoretical results, we first note that if we take the limit of $\kappa\to\infty$ and $\alpha\to\infty$ in Result 1 (in either order), the ICL generalization error reduces to the generalization error of simple ridgeless interpolation with isotropic Gaussian covariates in $d^{2}$ dimensions [36, 38]:
$\lim_{\alpha\to\infty}\lim_{\kappa\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}=\lim_{\kappa\to\infty}\lim_{\alpha\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}=\begin{dcases}1-\tau+\frac{\rho}{1-\tau}&\tau<1,\\ \frac{\rho\tau}{\tau-1}&\tau>1.\end{dcases}$ (33)
This limiting result makes sense, given that in this limit the ICL generalization problem reduces to the generalization error of ridge regression in $d^{2}$ dimensions with covariates formed as the tensor product of i.i.d. Gaussian vectors, which by universality results in [37] should in turn be asymptotically equal to that for isotropic Gaussian covariates [36].
3.3 Sample-wise double-descent
(a) Ridgeless ICL error against $\tau$
(b) Ridgeless IDG error against $\tau$
(c) Finite ridge IDG error against $\tau$
Figure 1: ICL performance as a function of $\tau$: theory (solid lines) vs simulations (dots). Plots of (a), (c) ICL error $e^{\text{ICL}}(\Gamma^{\ast})$ and (b) IDG error $e^{\text{IDG}}(\Gamma^{\ast})$ vs $\tau$ at optimal $\Gamma^{\ast}$. Parameters: $d=80,\kappa=0.5$, and $\rho=0.01$. Averages and standard deviations are computed over 10 runs.
How large should $n$, the pretraining dataset size, be for the linear attention to succesfully learn the ICL and IDG tasks? In Figure 1, we plot our theoretical predictions for the ICL and IDG generalization error as a function of $\tau=n/d^{2}$ and verify them with numerical simulations. Our results demonstrate that the quadratic scaling of sample size with input dimensions is indeed the appropriate regime for nontrivial learning phenomena to occur.
@@ 3.3 Sample-wise double-descent
3.4 The ICL error can have non-monotonic dependence on context length
(a) ICL error against $\alpha$
(b) IDG error against $\alpha$
@@ 3.4 The ICL error can have non-monotonic dependence on context length
Figure 2: ICL performance as a function of $\alpha$: theory (solid lines) vs simulations (dots). Plots of (a) ICL error $e_{g}^{\text{ICL}}(\Gamma^{\ast})$ and (b) IDG error $e_{g}^{\text{IDG}}(\Gamma^{\ast})$ vs $\alpha$ at optimal $\Gamma^{\ast}$. We highlight that, while the IDG error is monotonic in $\alpha$, the ICL error for $\kappa=0.5$ and $\kappa=0.75$ is non-monotonic. Parameters: $d=70,\tau=20,\rho=0.01$, ridgeless. Averages and standard deviations are computed over 10 runs.
@@
How large should the context length be? In Figure 2, we plot our theoretical results verified with experiments. We observe that we have correctly identified the regime where ICL appears: context length scales linearly with input dimensions. An interesting observation is that the ICL error does not always monotonically decrease with context length. There are parameter configurations with $\kappa<1$ (blue curve in Figure 2(a)) for which the ICL error is minimal at some finite $\alpha$.
3.5 Memorization to ICL transition with increasing pretraining task diversity
Recall that the parameter $\kappa=k/d$ controls the diversity of the training task vectors. How large should it be for ICL to emerge? Our theory corroborates a phenomenon that was empirically observed in a recent study [18]. Figure 3 shows a transition in the nature of the predictions that the linear attention module makes. For low $\kappa$, the model’s performance is close to that of the dMMSE estimator. This indicates that the model inherently assumes the task vector is one of the $k$ vectors encountered in its pretraining dataset, effectively memorizing these task vectors. As $\kappa$ increases beyond 1, the model’s performance approaches that of the ridge estimator. In this regime, the model generalizes to task vectors beyond its pretraining dataset, behaving as if it has learned the true prior on the task vectors despite having only seen a finite subset in the pretraining dataset.
Figure 3: Comparison of linear Transformer generalization error with the dMMSE estimator given by eq. 24: theory (solid lines) vs simulations (dots, triangles). Each value plotted is the excess value of generalisation error over the noise level $\rho$. Parameters: $\tau=0.2\alpha,\rho=0.01$, ridgeless. Averages and standard deviations for linear model are computed over 10 runs.
To further understand the role of $\kappa$ in the solution learned by the linear attention mechanism, consider the regime where $\tau,\alpha\to\infty$ with $\tau/\alpha=c^{\ast}$ kept fixed. Under this setting, we have

### v2

@@ II.5 Model pretraining
We estimate model parameters by minimizing MSE loss on next-output prediction with ridge regularistion, giving
@@
where $\lambda>0$ is a regularization parameter, and $H_{Z^{\mu}}$ refers to the input matrix (17) populated with the $\mu$th sample sequence. The factor $n/d$ in front of $\lambda$ makes sure that, when we take the $d\to\infty$ or $n\to\infty$ limits later, there is still a meaningful ridge regularization when $\lambda>0$. The solution to the optimization problem in (20) can be expressed explicitly as
$\text{vec}(\Gamma^{\ast})=\left(\frac{n}{d}\lambda I+\sum_{\mu=1}^{n}\text{vec}(H_{Z^{\mu}})\text{vec}(H_{Z^{\mu}})^{\top}\right)^{-1}\sum_{\mu=1}^{n}y_{\ell+1}^{\mu}\text{vec}(H_{Z^{\mu}})$ (21)
@@ II.6 Evaluation
$e(\Gamma)\coloneqq\mathbb{E}_{\mathcal{P}_{\mathrm{test}}}\left[\left(y_{\ell+1}-\langle\Gamma,H_{Z}\rangle\right)^{2}\right],$ (22)
where $(Z,y_{\ell+1})\sim\mathcal{P}_{\mathrm{test}}$ is a new sample drawn from the probability distribution of the test dataset. We consider two different test data distributions $\mathcal{P}_{\mathrm{test}}$:
@@
    ICL task: $x_{i}$ and $\epsilon_{i}$ are i.i.d. Gaussians as in the pretraining case. However, each task vector $w^{\text{test}}$ associated with a test input sequence of length $\ell$ is drawn independently from $\mathcal{N}(0,I_{d})$. We will denote the test error under this setting by $e^{\mathrm{ICL}}(\Gamma)$.
@@
    In-distribution generalization (IDG) task: The test data are generated in exactly the same manner as the training data, i.e., $\mathcal{P}_{\mathrm{test}}=\mathcal{P}_{\mathrm{train}}$, hence the term in-distribution generalization. In particular, the set of unique task vectors $\{w_{1},\ldots,w_{k}\}$ is identical to that used in the pretraining data. We will denote the test error under this setting by $e^{\text{IDG}}(\Gamma)$. This task can also be referred to as in-weight learning.
The ICL task evaluates the true in-context learning performance of the linear attention module. The task vectors in the test set differ from those seen in training, requiring the model to infer them from context. The IDG task assesses the model’s performance on task vectors it has previously encountered during pretraining. High performance on the IDG task but low performance on the ICL task indicates that the model memorizes the training task vectors. Conversely, high performance on the ICL task suggests that the model can learn genuinely new task vectors from the provided context.
To understand the performance of our model on both ICL and IDG tasks, we will need to evaluate these expressions for the pretrained attention matrix $\Gamma^{\ast}$ given in (21). An asymptotically precise prediction of $e^{\mathrm{ICL}}(\Gamma^{\ast})$ and $e^{\mathrm{IDG}}(\Gamma^{\ast})$ will be a main result of this work.
II.7 Baselines: Bayes-optimal within-context estimators
Following [20], it is useful to compare the predictions made by the trained linear attention to optimal estimators that use only the current context information. These estimators rely solely on the data within the given context for their predictions, but need oracle knowledge of the full statistical models underlying the data. Under the mean square loss, the optimal Bayesian estimator $\hat{y}_{\text{Bayes}}=\mathbb{E}_{\mathcal{P}_{\mathrm{test}}}[y_{\ell+1}|x_{1},y_{1},x_{2},y_{2},\ldots,x_{\ell},y_{\ell},x_{\ell+1}]$ in our setting has the form
$\hat{y}_{\text{Bayes}}=(w_{\text{Bayes}})^{\top}x_{\ell+1},$ (23)
where $w_{\text{Bayes}}$ is the Bayes estimator of the task vector $w$.
@@ II.7 Baselines: Bayes-optimal within-context estimators
$w_{\text{Bayes}}^{\text{ridge}}\coloneqq\left(\sum_{i=1}^{\ell}x_{i}x_{i}^{\top}+\rho I_{d}\right)^{-1}\left(\sum_{i=1}^{\ell}y_{i}x_{i}\right),$ (24)
where the ridge parameter is set to the noise variance $\rho$. We will refer to it as the ridge estimator. For the IDG task, the Bayes-optimal estimator is given by
$w_{\text{Bayes}}^{\text{dMMSE}}\coloneqq\frac{\sum_{j=1}^{k}w_{j}e^{-\frac{1}{2\rho}\sum_{i=1}^{\ell}\left(y_{i}-w_{j}^{\top}x_{i}\right)^{2}}}{\sum_{j=1}^{k}e^{-\frac{1}{2\rho}\sum_{i=1}^{\ell}\left(y_{i}-w_{j}^{\top}x_{i}\right)^{2}}}.$ (25)
@@
$e^{\text{Bayes}}_{\mathcal{P}_{\mathrm{test}}}=\mathbb{E}_{\mathcal{P}_{\mathrm{test}}}\left[\left(y_{\ell+1}-(w_{\text{Bayes}})^{\top}x_{\ell+1}\right)^{2}\right],$ (26)
where $\mathcal{P}_{\mathrm{test}}$ can be associated with either the ICL or the IDG task, and $w_{\text{Bayes}}$ can be the ridge or the dMMSE estimator. To avoid possible confusion, we emphasize that we will sometimes plot the performance of an estimator on a task for which it is not optimal. For example, we will test the dMMSE estimator, which is Bayes-optimal for the pretraining distribution, on the ICL task, where it is not optimal. This will be done for benchmarking purposes.
III Theoretical results
To answer the questions raised in the introduction, we provide a precise asymptotic analysis of the learning curves of the linear attention module for ICL of linear regression. We then focus on various implications of these equations, and verify through simulations that our insights gained from this theoretical analysis extend to more realistic nonlinear Transformers.
III.1 Joint asymptotic limit
We have now defined both the structure of the training data as well as the parameters to be optimized. For our theoretical analysis, we consider a joint asymptotic limit in which the input dimension $d$, the pretraining dataset size $n$, the context length $\ell$, and the number of task vectors in the training set $k$, go to infinity together such that
@@ III.1 Joint asymptotic limit
The intuition for these scaling parameters can be seen as follows. Standard results in linear regression [36, 37, 38] show that to estimate a $d$-dimensional task vector $w$ from the $\ell$ samples within a context, one needs at least $\ell=\Theta(d)$. The number of unique task vectors that must be seen to estimate the covariance matrix of the true $d$-dimensional task distribution $\mathcal{N}(0,I_{d})$ should also scale with $d$, i.e. $k=\Theta(d)$. Finally, we see from (16) that the number of linear attention parameters to be learned is $\Theta(d^{2})$. This suggests that the number of individual contexts the model sees during pretraining should scale similarly, i.e., $n=\Theta(d^{2})$.
III.2 Learning curves for ICL of linear regression by a linear attention module
Our theoretical analysis, explained in detail in the Supplementary Information, leads to asymptotically precise expressions for the generalization errors under the two test distributions being studied. Specifically, our theory predicts that, as $d,n,\ell,k\to\infty$ in the joint limit given in (27),
@@ III.2 Learning curves for ICL of linear regression by a linear attention module
where $e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$ and $e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)$ are two deterministic functions of the parameters $\tau$, $\alpha$, $\kappa$, $\rho$ and $\lambda$. The exact expressions of these two functions can be found in Section SI.5.2 and Section SI.5.3 of the SI, respectively. For simplicity, we only present in what follows the ridgeless limit (i.e., $\lambda\to 0^{+}$) of the asymptotic generalization errors.
@@
$q^{\ast}\coloneqq\frac{1+\rho}{\alpha},\quad m^{\ast}\coloneqq\mathcal{M}_{\kappa}\left({q^{\ast}}\right),\quad\mu^{\ast}\coloneqq q^{\ast}\mathcal{M}_{\kappa/\tau}(q^{\ast}),$ (30)
where $\mathcal{M}_{\kappa}(\cdot)$, defined in (SI.8.3), is a function related to the Stieltjes transform of the Marchenko-Pastur law. Then
@@
[…]
where $\mathcal{M}^{\prime}_{\kappa}(\cdot)$ means the derivative of $\mathcal{M}_{\kappa}(q)$ with respect to $q$.
@@
[…]
@@
We will discuss various implications of these equations in the next sections. We derived these results using techniques from random matrix theory. The full setup and technical details are presented in the Supplementary Information in Section SI.1 through Section SI.5. The computations involve analysis of the properties of the finite-sample optimal parameter matrix $\Gamma^{*}$ (see (21)). A key technical component of our analysis involves characterizing the spectral properties of the sample covariance matrix of $n=\Theta(d^{2})$ i.i.d. random vectors in dimension $\Theta(d^{2})$. Each of these vectors is constructed as the vectorized version of the matrix in (17). Related but simpler versions of this type of random matrices involving the tensor product of i.i.d. random vectors have been studied in recent work [39]. Some of our derivations are based on non-rigorous yet technically plausible heuristics. We support these predictions with numerical simulations and discuss in the Supplementary Information the steps required to achieve a fully rigorous proof.
III.3 Sample-wise double-descent
((a)) Ridgeless IDG error against $\tau$
((b)) Ridgeless ICL error against $\tau$
((c)) Finite ridge ICL error against $\tau$
Figure 1: ICL performance as a function of $\tau$: theory (solid lines) vs simulations (dots). Plots of (a) $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$, (b) $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$, and (c) $e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$ against $\tau$. Simulated errors calculated by evaluating the corresponding test error on the corresponding optimized $\Gamma^{*}$. Parameters: $d=100$, $\rho=0.01$ for all; (a), (b) $\kappa=0.5$, (c) $\alpha=10,\kappa=\infty$. Averages and standard deviations are computed over 10 runs.
How large should $n$, the pretraining dataset size, be for the linear attention to succesfully learn the task in-context? In Figure 1, we plot our theoretical predictions for the ICL and IDG error as a function of $\tau=n/d^{2}$ and verify them with numerical simulations. Our results demonstrate that the quadratic scaling of sample size with input dimensions is indeed an appropriate regime where nontrivial learning phenomena can be observed.
@@ III.3 Sample-wise double-descent
Finally, we note that if we take the limit of $\kappa\to\infty$ and $\alpha\to\infty$ in Result 1 (in either order), the ICL generalization error in the ridgeless case reduces to the generalization error of simple ridgeless interpolation with isotropic Gaussian covariates in $d^{2}$ dimensions [38, 41]:
$\lim_{\alpha\to\infty}\lim_{\kappa\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ $=\lim_{\kappa\to\infty}\lim_{\alpha\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ (33)
=$=\begin{dcases}1-\tau+\frac{\rho}{1-\tau}&\tau<1,\\ \frac{\rho\tau}{\tau-1}&\tau>1.\end{dcases}$ (34)
This result makes sense, given that in this limit the ICL generalization problem reduces to the generalization error of ridge regression in $d^{2}$ dimensions with covariates formed as the tensor product of i.i.d. Gaussian vectors, which by universality results in [39] should in turn be asymptotically equal to that for isotropic Gaussian covariates [38]. We note that taking the limit $\alpha\to\infty$ is not strictly necessary for this universality, but doing so simplifies the expressions and makes the double-descent obvious.
III.4 ICL and IDG error curves can have non-monotonic dependence on context length
((a)) $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ against $\alpha$; $\tau=0.5,\rho=0.5$
((b)) $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ against $\alpha$; $\tau=0.5,\rho=0.1$
((c)) $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ against $\alpha$; $\tau=20,\rho=0.5$
@@ III.4 ICL and IDG error curves can have non-monotonic dependence on context length
Figure 2: Error curves as functions of $\alpha$: theory (solid lines) vs simulations (dots). Plots of (a) $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ and (b), (c) $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ against $\alpha$. Hollowed markers, when plotted, indicate $\alpha^{\ast}$ value minimizing error if it exists. Parameters: $d=100$. Averages and standard deviations are computed over 20-100 runs.
((a)) Non-monotonicity of $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$, $\tau=0.5$
((b)) Non-monotonicity of $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$, $\tau=0.5$
((c)) Non-monotonicity of $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$, $\tau=20$
Figure 3: Phase diagram of non-monotonicity in $\alpha$ of ridgeless IDG and ICL error, for fixed $\tau$ against task diversity $\kappa$ and label noise $\rho$. The colormap corresponds to the value $\alpha^{\ast}$ that minimizes error in $\alpha$, if it exists, at that particular $\kappa,\rho$ pair; grey is plotted if the error curve is monotonic at that $\kappa,\rho$ pair. The dashed vertical lines are plotted at $\kappa=\text{min}(\tau,1)$. We know from (35) and surrounding discussion that ICL error diverges in $\alpha$ for all $\kappa,\rho$ to the left of this line, and IDG error diverges on this line for $\tau<1$. Letters correspond to matching parameter values of the curves plotted in their respective Figure 2 subfigure.
@@
How large should the context length be? In Figure 2, we plot our theoretical results verified with experiments. We observe that we have correctly identified the regime where in-context and in-weight learning appear: context length indeed scales linearly with input dimensions, as numerical simulations computed using finite $d$ fit the asymptotic error curves.
An interesting observation is that the IDG and ICL errors do not always monotonically decrease with context length; we see that there are parameter configurations for which the error curves are minimized at some finite $\alpha^{\ast}$.
While the functional form of $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ and $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ is too complex to study their minimizers analytically, we can investigate the non-monotonicity of IDG and ICL error in $\alpha$ numerically. This is done in Figure 3, where the blue-green colormap corresponds to the value $\alpha^{\ast}$, if it exists, that minimizes $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$ or $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$ at fixed $\tau,\kappa,\rho$. The grey color indicates cases where the corresponding function is monotonic in $\alpha$. The results reveal that the non-monotonicity of the errors with respect to context length $\alpha$ depends in a nontrivial manner on both the task diversity $\kappa$ and the noise-to-signal ratio $\rho$.
In addition to the non-monotonicity of error curves in $\alpha$, we also notice a divergence in error as $\alpha$ increases for some $\kappa,\tau$ values. To determine when this occurs, we compute the large $\alpha$ limit of the above error curves. For ICL error, we have